# Reexamining Paradigms of End-to-End Data Movement

### v4

@@ Abstract.
The pursuit of high-performance data transfer often focuses on raw network bandwidth, where international links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete. It equates provisioned link speeds with practical, sustainable data movement capabilities. It is a common observation that lower than desired data rates manifest even on 10 Gbps links and commodity hardware, with higher-speed networks only amplifying their visibility. We investigate six paradigms – from network latency and TCP congestion control to host-side factors such as CPU performance and virtualization – that critically impact data movement workflows. These paradigms represent widely “accepted” engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. We introduce the “Drainage Basin Pattern” conceptual model for reasoning about end-to-end data flow constraints across heterogeneous hardware and software components at varying desired data rates to address the fidelity gap between raw bandwidth and application-level throughput. Our findings are validated through rigorous production-scale deployments, from 10 Gbps links to U.S. DOE ESnet technical evaluations and transcontinental production trials over 100 Gbps operational links. The results demonstrate that principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design enables consistent, predictable performance for moving data at scale and speed.
@@ 1. Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is an unsolved foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps (Energy Sciences Network (2025a), ESnet). However, a predominant focus on high raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path— including storage, host architectures, software design, and security measures—rather than by the network alone. This holds true from 10 Gbps to 100 Gbps and faster.
@@
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II (Thayer and Fang, 2019), sustained petabyte transfers (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet), and international deployments (ICM, 2019)—our work demonstrates that once a link is in place, whether fast or high-speed, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations (Kissel and Fang, 2020c, b, a), historical transcontinental trials over 100 Gbps operational links, and high-performance validations on the TaiWan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (2025), TWAREN). Together, this body of evidence confirms that the bottlenecks have shifted from the wide-area fabric to the internal systems architecture of the endpoints.
A key architectural component for achieving predictable performance is the use of high-speed burst buffers (Bhimji et al., 2016) as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. Our prior work, including the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) (NSCC, 2019) and official ESnet evaluations (Kissel and Fang, 2020c, b, a), underscores that these factors, not the network, set the true performance limits of data movement, regardless of the network’s nominal speed.
@@ 2. Background / Related Work
Research programs such as the Global Research Platform (GRP) ((2025a), GRP) have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop ((2025b), GRP) and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective by examining data movement as a first-class systems problem (Fang et al., 2016; Fang, 2019).
@@ 2.1. Burst Buffers and Data Staging
In this architecture, the burst buffer serves both as a fast storage tier and as a deliberate decoupling mechanism. It isolates the data mover from production storage services, which are often optimized for capacity or ease of use, rather than throughput or latency. This decoupling allows each layer to operate within its effective performance regime while staging coordinates data movement across mismatched behaviors.
@@ 2.3. Integrated Appliances vs Software-Centric Approaches
The network-centric approach discussed previously, while valuable, does not fully address the processes required for achieving target data rates reliably in production. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; ICM, 2019; NSCC, 2019)—can create significant operational challenges. Integration, tuning, and maintenance require deep, multi-disciplinary expertise, presenting a substantial total cost of ownership that many organizations are not equipped to sustain.
@@
The resulting variability makes performance unpredictable and hard to reproduce across different organizations and environments.
@@
This paper demonstrates that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. This work builds on extensive prior advances in high-speed networking and host tuning, and focuses on architectural principles that make such expertise deployable and repeatable across production environments. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how to do it right. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
@@
The overall approach is grounded in technology developed over the past decade to meet the ambitious data movement requirements of the premier United States (U.S.) Department of Energy (DOE) Exascale Computing Preparation Project, LCLS-II (Thayer and Fang, 2019). The core challenges and performance targets defined by the project directly motivated the architecture. Furthermore, the ongoing dialogue with LCLS-II and other tier-1 scientific entities continues to advance and validate the evolution of this holistic approach to data movement systems. The core software’s performance has been independently validated on long-distance 100 Gbps links, encompassing petabyte-scale transfers with ESnet (Energy Sciences Network (2017), ESnet), (Energy Sciences Network (2018), ESnet), and (Kissel and Fang, 2020c), SCA19 DMC (NSCC, 2019), and transcontinental production trials (ICM, 2019).
@@
The outcomes documented by independent institutions, e.g., (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; NSCC, 2019) are reproducible; nevertheless, our experience shows that operational reproducibility depends critically on the availability of pre-validated system configurations rather than ad-hoc assembly at each site. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based units from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to micro servers (Fig. 3). This demonstrates that reliable, efficient data movement is most effectively achieved through complete systems engineering (Kissel and Fang, 2020b, a). The underlying principle is akin to a high-performance electric vehicle (EV): engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks. Just as an EV is not merely a carriage with an electric motor swapped in, a purposely-built data movement appliance is not simply a server with powerful CPUs, NVMe SSDs, and high-bandwidth NICs; the hardware Bill of Materials (BOM) is meticulously formed to maximize operational and cost efficiency. The components are orchestrated via software to eliminate the non-deterministic bottlenecks inherent in general-purpose computing. In passing we will note that attempts to introduce legacy, non-optimized software data movers on such a deterministic platform result in performance regressions, negating the advantages of the co-design principle.
@@ 3.1. The Latency Killer Paradigm
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base (Energy Sciences Network (2025b), ESnet) have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section 3.3. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table 2 for Ethernet adapter ring buffer settings and Table 3 for tuned Linux kernel parameters.
The test utilized iperf3, a benchmark tool maintained by the U.S. DOE ESnet (Energy Sciences Network (2025g), ESnet). Beginning with version 3.16 (released November 30, 2023 (Energy Sciences Network (2023), ESnet)), the tool was re-architected to be genuinely multithreaded, making it significantly easier to tune for high-latency, high-speed links (e.g., ~160 Gbps memory-to-memory over a 200 Gbps ESnet testbed data path) (Energy Sciences Network (2025d), ESnet).
@@
Table 3. Tuned Linux kernel parameters. This specific set was empirically determined to be optimal for the zx data mover’s concurrency model, whereas other data movers may require more extensive tuning. For example, see ESnet fasterdata, Linux Tuning (Energy Sciences Network (2025e), ESnet).
@@ 3.2.1. The Packet Loss Paradigm
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in (Energy Sciences Network (2009), ESnet)) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to 10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
@@
The operational lower bound on error rates is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of (Kissel and Fang, 2020c) (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in (Energy Sciences Network (2018), ESnet). Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This result directly contradicts the notion that packet loss is the dominant limiter in modern, well-engineered wide-area backbones.

### v5

@@ Abstract.
The pursuit of high-performance data transfer often focuses on raw network bandwidth. International links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete. It equates provisioned link speeds with practical, sustainable data movement capabilities. It is a common observation that lower than desired data rates manifest even on 10 Gbps links, with higher-speed networks only amplifying their visibility. We investigate six paradigms – from network latency and TCP congestion control to host-side factors such as CPU performance and virtualization – that critically impact data movement workflows. These paradigms represent widely “accepted” engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. We introduce the “Drainage Basin Pattern” conceptual model for reasoning about end-to-end data flow constraints across heterogeneous hardware and software components at varying desired data rates to address the fidelity gap between raw bandwidth and application-level throughput. Our findings are validated through rigorous production-scale deployments, from 10 Gbps links to U.S. DOE ESnet technical evaluations and transcontinental production trials over 100 Gbps operational links. The results demonstrate that principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design enables consistent, predictable performance for moving data at scale and speed.
@@ 1. Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is an unsolved foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps (Energy Sciences Network (ESnet), 2025a). However, a predominant focus on high raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path— including storage, host architectures, software design, and security measures—rather than by the network alone. This holds true from 10 Gbps to 100 Gbps and faster.
@@
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II (Thayer and Fang, 2019), sustained petabyte transfers (Energy Sciences Network (ESnet), 2017, 2018), and international deployments (ICM, 2019)—our work demonstrates that once a link is in place, whether fast or high-speed, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations (Kissel and Fang, 2020c, b, a), historical transcontinental trials over 100 Gbps operational links, and high-performance validations on the TaiWan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (TWAREN), 2025). Together, this body of evidence confirms that the bottlenecks have shifted from the wide-area fabric to the internal systems architecture of the endpoints.
A key architectural component for achieving predictable performance is the use of high-speed burst buffers (Bhimji et al., 2016) as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. As the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) (NSCC, 2019), our co-design principles have been validated in a highly competitive, multi-site trial spanning global infrastructure (Lee, 2023) (see slides 5 and 8-9, which detail the global topology and list Zettar Inc. as the winner among seven international teams). The 2020 official ESnet evaluations (Kissel and Fang, 2020c, b, a) further underscore that these factors, not the network, set the true performance limits of data movement, regardless of the network’s nominal speed.
A central insight emerging from this body of evidence is that co-design yields not only high peak performance but also crucial operational benefits. As we detail in Section 2.3, the ESnet evaluation found that a properly co-designed system can sustain consistent throughput across file sizes from 1 MiB to 1 TiB with a single configuration. This stands in contrast to software-centric approaches that often require file-size-range-based optimization. Such operational simplicity is not merely a convenience; it is a direct consequence of architectural co-design and a prerequisite for deployment in production environments where datasets and workflows vary widely.
@@ 2. Background / Related Work
Research programs such as the Global Research Platform (GRP) ((GRP), 2025b) have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop ((GRP), 2025a) and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective by examining data movement as a first-class systems problem (Fang et al., 2016; Fang, 2019).
@@ 2.1. Burst Buffers and Data Staging
In this architecture, the burst buffer serves both as a fast storage tier and as a deliberate decoupling mechanism. It isolates the data mover from production storage services, which are often optimized for capacity or ease of use, rather than throughput or latency. This decoupling allows each layer to operate within its effective performance regime while staging coordinates data movement across mismatched behaviors. In other words, it acts as a low-jitter Interface that buffers the stochastic throughput and latency of the non-deterministic source (production storage) to ensure a deterministic, high-bandwidth supply to the high-speed sink.
@@ 2.3. Integrated Appliances vs Software-Centric Approaches
The network-centric approach discussed previously, while valuable, does not fully address the processes required for achieving target data rates reliably in production. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials (Energy Sciences Network (ESnet), 2017, 2018; ICM, 2019; NSCC, 2019)—can create significant operational challenges. Hardware selection, system integration, tuning, and maintenance require deep, multi-disciplinary expertise. In many cases, hardware is also significantly overprovisioned, with enterprise servers deployed for workloads that could be served efficiently by compact and far less expensive systems. This combination presents a substantial total cost of ownership that is not affordable to many organizations.
@@
Furthermore, the resulting variability makes performance unpredictable and hard to reproduce across different organizations and environments.
In addition, the distinction between integrated appliances and software-centric approaches manifests not only in peak performance but also in operational complexity and configuration overhead. A co-designed system, in which hardware and software are engineered together, can achieve robust performance across diverse workloads with minimal tuning. In contrast, software-centric solutions often require per-dataset optimization, making them fragile and labor-intensive in production environments with heterogeneous data.
This difference is clearly illustrated by an evaluation conducted by ESnet, reported in a 2020 technical report (Kissel and Fang, 2020c). The authors compared the tuning requirements of a co-designed data mover with those of a representative commercial software-centric solution. Section 10 of the report, Lessons Learned, are particularly instructive:
“With a well-tuned infrastructure, it is feasible for a software data mover to use a single setting for a wide range of file sizes and various size distributions (i.e., file size histograms).”
“Although tuning takes the adjustments of only a few parameters and is straightforward, the results from properly and insufficiently tuned can be very significant (e.g., $\geq 200\%$).”
Performance stability under realistic operating conditions is therefore as important as peak throughput. Systems that achieve good benchmark results only under carefully tuned parameters and selected datasets place a significant burden on operators. They may lack the time or expertise required for continuous performance tuning. In large infrastructures—where datasets, workflows, and file-size distributions vary widely—such fragility directly translates into operational inefficiency and reduced productivity. By contrast, systems designed through integrated hardware–software co-design can deliver consistently high performance across diverse workloads with minimal configuration effort, making them substantially more suitable for production environments.
The report further notes (see page 9, Fig. 1) that achieving comparable results with a typical commercial software-centric solution required tuning based on the specific data-size range. In contrast, the evaluated co-designed system delivered consistent performance with a single configuration across file sizes ranging from 1 KiB to 1 TiB.
Such a single, broadly effective configuration—referred to as global tuning—does not imply rigidity. A well-architected system can support hierarchical tuning, in which global settings provide sensible defaults for most workloads, while per-task parameters—accessible through an intuitive interface such as a built-in WebUI—allow advanced users to override defaults for specialized transfers. These per-task settings take precedence, providing flexibility without sacrificing ease of use.
@@
This paper explains that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. This work builds on extensive prior advances in high-speed networking and host tuning, and focuses on architectural principles that make such expertise deployable and repeatable across production environments. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how to do it right. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
@@
The overall approach is grounded in technology developed over the past decade to meet the ambitious data movement requirements of the premier United States (U.S.) Department of Energy (DOE) Exascale Computing Preparation Project, LCLS-II (Thayer and Fang, 2019). The core challenges and performance targets defined by the project directly motivated the architecture. Furthermore, the ongoing dialogue with LCLS-II and other tier-1 scientific entities continues to advance and validate the evolution of this holistic approach to data movement systems. The core software’s performance has been independently validated on long-distance 100 Gbps links, encompassing petabyte-scale transfers with ESnet (Energy Sciences Network (ESnet), 2017), (Energy Sciences Network (ESnet), 2018), and (Kissel and Fang, 2020c), SCA19 DMC (NSCC, 2019), and transcontinental production trials (ICM, 2019). See also Figs. LABEL:fig:fig5,-6.
@@
The outcomes documented by independent institutions, e.g., (Energy Sciences Network (ESnet), 2017, 2018; NSCC, 2019) are reproducible; nevertheless, our experience shows that operational reproducibility depends critically on the availability of pre-validated system configurations rather than ad-hoc assembly at each site. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based units from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to micro servers (Fig. 3). The resulting appliances are simpler, faster, cleaner, and already validated, while also reducing deployment cost through the use of standard hardware and software components.
It is important to distinguish the data movement appliances discussed in this paper from typical enterprise appliances—a firewall, switches, or NAS. They manage a single stack: security, network, storage, often atop a proprietary operating system. In contrast, a data movement appliance is fundamentally more demanding to engineer. It must simultaneously orchestrate storage I/O, compute resources, high-speed networking, and security, all while executing a parallel, concurrent software data mover that sustains near-line-rate throughput across these interacting subsystems. This complexity places such appliances outside the domain of general-purpose server deployments.
The co-design principle addresses this challenge by enabling the use of regular Commercial Off-The-Shelf (COTS) hardware and standard software components, yielding a solution that is both cost-effective and operationally more manageable—even though the proprietary expertise required to achieve this integration resides with the appliance designers. This stands in contrast to two common but misguided assumptions: that one can simply take the software, install it on arbitrary hardware, and expect identical results; or that enterprise IT teams can treat these appliances as just another compute server to be assembled and tuned locally. Both approaches overlook the deterministic, cross-layer optimization that distinguishes a true data movement appliance from a general-purpose system running data movement software.
Reliable, efficient data movement is achieved most effectively through complete systems engineering (Kissel and Fang, 2020b, a). The underlying principle is akin to a high-performance electric vehicle (EV): engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks. Just as an EV is not merely a carriage with an electric motor swapped in, a purposely-built data movement appliance is not simply a server with powerful CPUs, NVMe SSDs, and high-bandwidth NICs; the hardware Bill of Materials (BOM) is meticulously formed to maximize operational and cost efficiency. The components are orchestrated via software to eliminate the non-deterministic bottlenecks inherent in general-purpose computing. In passing, we note that attempts to introduce legacy, non-optimized software data movers on such a deterministic platform result in performance regressions, negating the advantages of the co-design principle.
@@ 3.1. The Latency Killer Paradigm
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base (Energy Sciences Network (ESnet), 2025b) have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section 3.3. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table 2 for Ethernet adapter ring buffer settings and Table 3 for tuned Linux kernel parameters.
The test utilized iperf3, a benchmark tool maintained by the U.S. DOE ESnet (Energy Sciences Network (ESnet), 2025e). Beginning with version 3.16 (released November 30, 2023 (Energy Sciences Network (ESnet), 2023)), the tool was re-architected to be genuinely multithreaded, making it significantly easier to tune for high-latency, high-speed links (e.g., ~160 Gbps memory-to-memory over a 200 Gbps ESnet testbed data path) (Energy Sciences Network (ESnet), 2025d).
@@
Table 3. Tuned Linux kernel parameters. This specific set was empirically determined to be optimal for the zx data mover’s concurrency model, whereas other data movers may require more extensive tuning. For example, see ESnet fasterdata, Linux Tuning (Energy Sciences Network (ESnet), 2025f).
@@ 3.2.1. The Packet Loss Paradigm
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in (Energy Sciences Network (ESnet), 2009)) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to 10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
@@
The operational lower bound on error rates is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of (Kissel and Fang, 2020c) (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in (Energy Sciences Network (ESnet), 2018). Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This result directly contradicts the notion that packet loss is the dominant limiter in modern, well-engineered wide-area backbones.